OpenStudy (barrelracer011):

Can someone explain to me how to do negative exponents... for example 2^-3

OpenStudy (ipwnbunnies):

\[a ^{-b} = \frac{ 1 }{ a^{b} }\]

OpenStudy (barrelracer011):

What?

OpenStudy (ipwnbunnies):

2^(-3) is the same as 1/(2^3)...

OpenStudy (ipwnbunnies):

Or 1/(2x2x2)

OpenStudy (barrelracer011):

I still dont get what your saying

OpenStudy (ipwnbunnies):

Do you know what 2^3 is.

OpenStudy (barrelracer011):

Haha I got it now thanks :) its just dividing that number as many times as the exponet

OpenStudy (mathstudent55):

Rule: \(a^{-b} = \dfrac{1}{a^b} \) Examples: \( 5^{-1} = \dfrac{1}{5^1} = \dfrac{1}{5} \) \(4^{-2} = \dfrac{1}{4^2} = \dfrac{1}{16} \) \(6^{-5} = \dfrac{1}{6^5} \)

OpenStudy (mathstudent55):

This is one way of thinking of negative exponents. In this example I started with \(5^4\). After the first line, each line is 5 times smaller that the line above. \(5^4 = 5 \times 5 \times 5 \times 5 = 625 \) \(5^3 = 5 \times 5 \times 5 = 125\) \(5^2 = 5 \times 5 = 25\) \(5^1 = 5 \) \(5^0 = 1 \) \(5^{-1} = \dfrac{1}{5} \) \(5^{-2} = \dfrac{1}{5 \times 5}=\dfrac{1}{25} \) \(5^{-3} = \dfrac{1}{5 \times 5 \times 5}=\dfrac{1}{125} \) Etc.